# Diffusion Generative Flow Samplers: Improving learning signals through partial trajectory optimization

### v1

@@ Abstract
We tackle the problem of sampling from intractable high-dimensional density functions, a fundamental task that often appears in machine learning and statistics. We extend recent sampling-based approaches that leverage controlled stochastic processes to model approximate samples from these target densities. The main drawback of these approaches is that the training objective requires full trajectories to compute, resulting in sluggish credit assignment issues due to use of entire trajectories and a learning signal present only at the terminal time. In this work, we present Diffusion Generative Flow Samplers (DGFS), a sampling-based framework where the learning process can be tractably broken down into short partial trajectory segments, via parameterizing an additional “flow function”. Our method takes inspiration from the theory developed for generative flow networks (GFlowNets), allowing us to make use of intermediate learning signals and benefit from off-policy exploration capabilities. Through a variety of challenging experiments, we demonstrate that DGFS results in more accurate estimates of the normalization constant than closely-related prior methods.
@@ 1 Introduction
In this work, we build on top of recent methods (Zhang & Chen, 2022; Vargas et al., 2023a) that train diffusion models for sampling, and reinterpret them from a GFlowNet perspective. With this in mind, we then propose the Diffusion Generative Flow Sampler (DGFS), a enhanced training method of diffusion models for sampling. In particular, we formulate a training approach that learns intermediate “flow [^1] In this work, the term “flow” specifically denotes the flow network structure in GFlowNets, which has nothing to do with normalizing flow models. functions” that aggregate information for training intermediate states. This allows us to further define partial trajectory-based training objectives, where we have access to only partial specification of the sampling path. We find that this approach significantly reduces gradient noise and helps stabilize convergence. Furthermore, our training framework enables us to receive learning signals at intermediate steps, even before the sequential sampling is completed, in contrast to previous methods that can only receive signals at the terminal steps. We provide our open-source PyTorch implementation at https://github.com/zdhNarsil/Diffusion-Generative-Flow-Samplers.
@@ 2.1 Sampling as stochastic optimal control
We aim to sample from a $D$-dimensional target distribution $\pi(\cdot)$ in $\mathbb{R}^{D}$ defined by a given unnormalized target density $\mu(\cdot)$ where $\pi(\cdot)=\frac{\mu(\cdot)}{Z}$ and $Z=\int\mu(\mathbf{x})\mathop{}\!\mathrm{d}\mathbf{x}$. Following a similar setup to prior works such as Zhang & Chen (2022) and Vargas et al. (2023a), we consider a sequential latent variable model with a joint distribution denoted $\mathcal{Q}$ of which samples follow a Markov process,
@@
Through this sequence model, we use the marginal distribution $\mathcal{Q}(\mathbf{x}_{N})$ at the terminal step $N$ to approximate $\pi(\mathbf{x}_{N})$. Here $p^{\text{ref}}_{0}(\cdot)$ is a reference distribution at time $n=0$, and $P_{F}(\cdot|\cdot)$ denotes a forward transition probability (or policy) distribution. This forward policy depends on $n$, i.e. it is $P_{F}^{n+1|n}(\cdot|\mathbf{x}_{n})$, but we omit the step index $n$ out of simplicity or assume $n$ is encoded in $\mathbf{x}_{n}$. We thus refer to $\mathcal{Q}$ as the forward process in the following text [^2] Note that this is consistent with prior GFlowNet literature, though it is the opposite direction considered by standard diffusion models in generative modeling, where “forward process” is colloquially used to denote the process from clean sample to white noise.. In this work, we constrain the transition to be a unimodal Gaussian distribution:
@@
It can be shown that the marginal distribution of $\mathcal{P}$ at the terminal time $N$ is exactly proportional to the target $\mu(\cdot)$. Hence it is natural to consider learning the drift $\mathbf{f}(\cdot,\cdot)$ in Equation 2 by minimizing the KL divergence $\mathcal{D}_{\mathrm{KL}}\left(\mathcal{Q}\|\mathcal{P}\right)$ between the forward process and this target process, which is equivalent to the following optimal control problem:
@@
If we keep the total time $T=hN$ fixed and push the number of steps $N\to\infty$, this formulation recovers the continuous-time stochastic optimal control formulation seen in prior works:
@@
where the forward process is defined by a variance exploding stochastic differential equation (VE-SDE), and $\mathop{}\!\mathrm{d}\mathbf{W}_{t}$ is a standard Wiener process. This formulation appears in standard stochastic optimal control theory (Kappen, 2005). The seminal path integral sampler (PIS) (Zhang & Chen, 2022) proposes to solve such optimal control problems with deep learning to tackle the sampling problem. Its follow-up, the denoising diffusion sampler (DDS) (Vargas et al., 2023a), takes a similar approach but adopts a variance preserving (VP) SDE rather than the variance exploding ones in Equation 5 and 7. We defer the derivation of VP modeling to Appendix B.2. The practical implementations of these prior works are performed in discrete time with constant step sizes, equivalent to the discrete formulation outlined above. Hence we also adopt the discrete-time formulation in the following sections for simplicity and better amenability to our proposed algorithm.
@@ 2.2 GFlowNets
Let $\mathcal{G}=(\mathcal{S},\mathcal{A})$ be a directed acyclic graph (DAG), where the set of vertices $\mathbf{s}\in\mathcal{S}$ are referred to as states, and the set of directed edges $(\mathbf{s}\to\mathbf{s}^{\prime})\in\mathcal{A}\subseteq\mathcal{S}\times\mathcal{S}$ represent the transitions between the states and are called actions. Given a designated initial state $\mathbf{s}_{0}$ with no incoming edges and a set of terminal states $s_{N}$ without outgoing edges, a complete trajectory can be defined as a sequence of states $\tau=(\mathbf{s}_{0}\to\mathbf{s}_{1}\to\ldots\to\mathbf{s}_{N})\in\mathcal{T}$. We define the forward policy $P_{F}(\mathbf{s}^{\prime}|\mathbf{s})$ as a distribution over the children of every non-terminal state $\mathbf{s}$, which induces a distribution over complete trajectories, $P_{F}(\tau)=\prod_{n=0}^{N-1}P_{F}(\mathbf{s}_{n+1}|\mathbf{s}_{n})$. This in turn defines a marginal distribution over terminating states $\mathbf{x}$ in the sense that $P_{T}(\mathbf{x})=\sum_{\tau\to\mathbf{x}}P_{F}(\tau)$, where we call $P_{T}(\cdot)$ the GFlowNet’s terminating distribution. The aim of GFlowNet learning is to obtain a policy such that $P_{T}(\cdot)\propto R(\cdot)$, where $R(\cdot):\mathcal{X}\to\mathbb{R}_{+}$ is a given reward function or unnormalized density, where we do not know the normalizing factor $Z=\sum_{\mathbf{x}}R(\mathbf{x})$. We also use the trajectory flow function $F(\tau)=ZP_{F}(\tau)$ to incorporate the effect of the partition function. Accordingly, the state flow function is the marginalization of the trajectory flow function $F(\mathbf{s})=\sum_{\tau\owns\mathbf{s}}F(\tau):\mathcal{S}\to\mathbb{R}_{+}$. Most previous GFlowNet works are about composite discrete objects such as molecules, with only Lahlou et al. (2023) providing theoretical support for GFlowNets on continuous spaces. Our work aims to tackle practical complex sampling problems in continuous domains, i.e. $\mathcal{S}\subseteq\mathbb{R}^{D}$.
@@ 3.2 Updating parameters with incomplete trajectories
This formulation is similar to the GFlowNet DB loss in Equation 9. Actually, it is not hard to see that the above modeling is a variant of GFlowNets. In the GFlowNet language, the target density $\mu(\cdot)$ is the terminal reward function, the temporal-spatial specification space $\{(n,\mathbf{x}_{n}):n\in\mathbb{N},\mathbf{x}_{n}\in\mathbb{R}^{D}\}$ is the GFlowNet state space $\mathcal{S}$, the forward transition probability $P_{F}(\cdot|\cdot)$ of the diffusion process $\mathcal{Q}$ is the GFlowNet forward policy, the backward transition probability $P_{B}(\cdot|\cdot)$ of the reference process $\mathcal{Q}^{\text{ref}}$ is the GFlowNet backward policy, and the helper network $F_{n}(\cdot)$ is the state flow $F(\mathbf{s})$. Notice here that the backward policy is fixed and does not contain learnable parameters. In our setting, we use the step index $n$ to distinguish between intermediate states $\{(n,\mathbf{x}_{n}):n<N,\mathbf{x}_{n}\in\mathbb{R}^{D}\}$ and terminating states $\{(N,\mathbf{x}_{N}),\mathbf{x}_{N}\in\mathbb{R}^{D}\}$ [^3] Without loss of generality, we slightly overuse the notations and simply use $\mathbf{x}_{n}$ to denote $(n,\mathbf{x}_{n})$.. What’s more, the goal of GFlowNets, which is to generate terminal states with probability proportional to the reward value, matches the goal of sampling in our setting.
@@
where $\lambda\in\mathbb{R}_{+}$ is a scalar for controlling assigned weights to different partial trajectories.
@@
Exploration with off-policy training.
The usage of the KL divergence in previous works restricts the training to be on-policy, as the training trajectory has to come from the model itself. Using off-policy samples is possible only with importance sampling, which typically results in high variance. This limits the exploration ability of these methods. On the other hand, notice that our objectives in Equation 9 or Equation 14 do not require the training samples to follow any particular distribution (only to have full support), which means our method supports off-policy training without importance sampling. Therefore, we employ off-policy exploration by using a larger variance coefficient in the forward process during the rollout stage. This helps DGFS to better capture diversity in complex multimodal target distributions. See more analysis in Appendix C.7.
@@
where $B\in\mathbb{N}_{+}$ is the number of particles (i.e., trajectories) used for the estimation. Similar to the importance weighted auto-encoders (Burda et al., 2016), our estimator provides a valid lower bound of the logarithm of the normalizing factor with any forward process $\mathcal{Q}$. In this work, we use the difference between estimated log partition function and the ground truth as an evaluation metric. See the Appendix C.2 for additional analysis.
@@ 3.3 Discussion
Figure 2 shows that the variance of stochastic gradients with DGFS is much smaller than that with PIS; notice that the comparison is valid as we take the same neural network architecture for both algorithms. We hypothesize that DGFS benefits from the low variance behavior due to the following two reasons: (1) From the perspective of temporal difference learning (Sutton & Barto, 2018; Kearns & Singh, 2000), it has long been discussed that policy learning from shorter trajectory data would lower the variance at the cost of potentially higher bias. Before this work, it was impossible for sampling with stochastic process algorithms to update with partial trajectories, while in the proposed DGFS method, we use an additional flow network to make this feasible. In Figure 4 we also show that the flow network can be properly learned and thus does not introduce much bias, indicating that DGFS achieves a better bias-variance trade-off. (2) Probabilistic ML researchers have shown that for KL divergence based distribution learning, the gradient will not vanish even if the optimal distribution has been reached; see Roeder et al. (2017) and Section 4 of Xu et al. (2022). On the other hand, the DGFS objective does achieve zero gradient with the optimal solution (c.f. Appendix C.3), making it a more stable algorithm.
@@
Previous theoretical studies (Bortoli, 2022; Chen et al., 2022; Lee et al., 2022) have demonstrated that the distribution of terminal samples generated by a diffusion model can converge to the target distribution under mild assumptions if the control term is well learned. These findings are also applicable to DGFS since these proofs are independent of how the networks are trained. Additionally, Zhang et al. (2023b) have proven that a perfectly learned score term corresponds to a zero GFlowNet training loss under mild conditions, following theory of GFlowNets (Bengio et al., 2021; 2023). This provides assurance that a sufficiently well-trained DGFS can accurately sample from the target distribution.
@@ 4 Related works
MCMC methods (Brooks et al., 2011) are the most classical sampling algorithms, with Langevin dynamics (Welling & Teh, 2011) and Hamilton Monte Carlo (Duane et al., 1987; Hoffman & Gelman, 2011; Chen et al., 2014) as representatives. With MCMC operators as building blocks, sequential MC methods (Liu, 2001; Doucet et al., 2001) generate samples ancestrally through a series of annealed targets. On the other hand, people turn to parametric models such as deep networks to learn samplers via variational inference (Rezende & Mohamed, 2015). One the most commonly adopted models is the normalizing flow (Dinh et al., 2014) due to its tractability and expressiveness. Various work have been conducted to improve normalizing flow based sampling (Müller et al., 2018; Gao et al., 2020; Wu et al., 2020; Chen et al., 2020; Gabri’e et al., 2021; Arbel et al., 2021; Midgley et al., 2022; de G. Matthews et al., 2022).
Recently, many researchers have developed powerful diffusion process to model distributions in the generative modeling community (Vincent, 2011; Sohl-Dickstein et al., 2015; Ho et al., 2020a; Song et al., 2020). This inspires people to study how to utilize diffusion process as samplers to learn distribution from given unnormalized target densities (Tzen & Raginsky, 2019; Zhang & Chen, 2022; Geffner & Domke, 2022; Doucet et al., 2022; Vargas et al., 2023a; Richter et al., 2023; Vargas & Nusken, 2023; Vargas et al., 2023b). Our work also falls into this category.
@@
GFlowNet (Bengio et al., 2023) is a family of generalized variational inference algorithms with a focus on treating the sampling process as a sequential decision-making process. It is original proposed to sample diversity-seeking objects in structured scientific domains such as biological seuqence and molecule design (Bengio et al., 2021; Jain et al., 2022; 2023b; 2023a; Shen et al., 2023; Zhang et al., 2023d; Hernández-García et al., 2023). However, people have developed GFlowNet theories and methodologies to be suitable for generating general objects with composite structures. GFlowNet has rich connection with existing probabilistic ML methods (Zhang et al., 2022b; Zimmermann et al., 2022; Malkin et al., 2023; Zhang et al., 2022a; Hu et al., 2023). As a policy learning framework under Markov decision processes (MDPs), researchers have also drawn fruitful connections between GFlowNet and reinforcement learning (Pan et al., 2023b; c; a). GFlowNet has also been demonstrated to have wide application in causal discovery (Deleu et al., 2022; 2023; Atanackovic et al., 2023), feature attribution (Li et al., 2023a), network structure learning (Liu et al., 2022), job scheduling (Zhang et al., 2023a), and graph combinatorial optimization (Zhang et al., 2023c).
Table 1: Experimental results on benchmarking target densities. For each algorithm, we report its absolute estimation bias for the log normalizing factor and the corresponding standard deviation.
@@
DDS $0.028$±$\pm 0.013$ $0.416$±$\pm 0.094$ $1.154$±$\pm 0.626$ $1.740$±$\pm 1.158$ N/A444 Due to time constraints, we are unable to reproduce DDS performance using PyTorch for this task. Instead, we encountered “not a number”, probably due to a mismatch in implementation details or other unspecified difference between PyTorch and Jax. The DDS results reported in its original paper can be found here.
DGFS $\bf{0.019}$±$\pm 0.008$ $\bf{0.274}$±$\pm 0.014$ $\bf{0.904}$±$\pm 0.067$ $\bf{0.180}$±$\pm 0.083$ $\bf{8.974}$±$\pm 1.169$
Most prior works focus on structured discrete data space. A prior study (Li et al., 2023b) initially explored GFlowNets in a continuous space, whereas their theory is demonstrated to have non-trivial flaws when considering integration under continuous measure as indicated by Lahlou et al. (2023). This work provides a practical and scalable algorithm for sampling problems in continuous spaces.

### v3

@@ Abstract
We tackle the problem of sampling from intractable high-dimensional density functions, a fundamental task that often appears in machine learning and statistics. We extend recent sampling-based approaches that leverage controlled stochastic processes to model approximate samples from these target densities. The main drawback of these approaches is that the training objective requires full trajectories to compute, resulting in sluggish credit assignment issues due to use of entire trajectories and a learning signal present only at the terminal time. In this work, we present Diffusion Generative Flow Samplers (DGFS), a sampling-based framework where the learning process can be tractably broken down into short partial trajectory segments, via parameterizing an additional “flow function”. Our method takes inspiration from the theory developed for generative flow networks (GFlowNets), allowing us to make use of intermediate learning signals. Through various challenging experiments, we demonstrate that DGFS achieves more accurate estimates of the normalization constant than closely-related prior methods.
@@ 1 Introduction
In this work, we build on top of Lahlou et al. (2023), who noted that recent methods that train diffusion models given an unnormalized density function (Zhang & Chen, 2022; Vargas et al., 2023a) can be interpreted from a GFlowNet perspective (Bengio et al., 2023). Although most GFlowNet works perform amortized inference on combinatorial discrete domains, previous work (Lahlou et al., 2023) establishes a theoretical foundation for continuous space GFlowNets and conducts numerical analysis to show their feasibility. Based on this framework, we propose the Diffusion Generative Flow Sampler (DGFS), an enhanced training method of diffusion models for sampling when given an unnormalized density function. In particular, we formulate a training approach that learns intermediate “flow [^1] In this work, the term “flow” specifically denotes the flow network structure in GFlowNets, which has nothing to do with normalizing flow models. functions” that aggregate information for training intermediate states. This allows us to further define partial trajectory-based training objectives, where we have access to only partial specification of the sampling path. We find that this approach significantly reduces gradient noise and helps stabilize convergence. Furthermore, our training framework enables us to receive learning signals at intermediate steps, even before the sequential sampling is completed, in contrast to previous methods that can only receive signals at the terminal steps.
@@ 2.1 Sampling as stochastic optimal control
We aim to sample from a $D$-dimensional target distribution $\pi(\cdot)$ in $\mathbb{R}^{D}$ defined by a given unnormalized target density $\mu(\cdot)$ where $\pi(\cdot)=\frac{\mu(\cdot)}{Z}$ and $Z=\int\mu(\mathbf{x})\mathop{}\!\mathrm{d}\mathbf{x}$. Following a similar setup to prior works such as Tzen & Raginsky (2019); Zhang & Chen (2022); Vargas et al. (2023a), we consider a sequential latent variable model with a joint distribution denoted $\mathcal{Q}$ of which samples follow a Markov process,
@@
Through this sequence model, we use the marginal distribution $\mathcal{Q}(\mathbf{x}_{N})$ at the terminal step $N$ to approximate $\pi(\mathbf{x}_{N})$. Here $p^{\text{ref}}_{0}(\cdot)$ is a reference distribution at time $n=0$, and $P_{F}(\cdot|\cdot)$ denotes a forward transition probability (or policy) distribution. This forward policy depends on $n$, i.e. it is $P_{F}^{n+1|n}(\cdot|\mathbf{x}_{n})$, but we omit the step index $n$ out of simplicity or assume $n$ is encoded in $\mathbf{x}_{n}$. We thus refer to $\mathcal{Q}$ as the forward process in the following text [^2] Note that this is consistent with prior GFlowNet literature, though it is the opposite direction considered by standard diffusion model literature in generative modeling, where “forward process” is colloquially used to denote the process from clean sample to white noise.. In this work, we constrain the transition to be a unimodal Gaussian distribution:
@@
It can be shown that the marginal distribution of $\mathcal{P}$ at the terminal time $N$ is exactly proportional to the target $\mu(\cdot)$ by marginalizing out the variables $\mathbf{x}_{0},\ldots,\mathbf{x}_{N-1}$ in $\mathcal{P}$; see Theorem 1 of Zhang & Chen (2022). Hence it is natural to consider learning the drift $\mathbf{f}(\cdot,\cdot)$ in Equation 2 by minimizing the KL divergence $\mathcal{D}_{\mathrm{KL}}\left(\mathcal{Q}\|\mathcal{P}\right)$ between the forward process and this target process, which is equivalent to the following optimal control problem:
@@
If we keep the total time $T=hN$ fixed and push the number of steps $N\to\infty$, this formulation recovers the continuous-time stochastic optimal control formulation seen in prior works (Zhang & Chen, 2022):
@@
where the forward process is defined by a variance exploding stochastic differential equation (VE-SDE) [^3] To avoid confusion, we would like to remark again that in literature the term “forward process” is usually used to describe the noising process, which is actually variance shrinking in PIS; see Equation 21., and $\mathop{}\!\mathrm{d}\mathbf{W}_{t}$ is a standard Wiener process. This formulation appears in standard stochastic optimal control theory (Kappen, 2005). The seminal path integral sampler (PIS) (Zhang & Chen, 2022) proposes to solve such optimal control problems with deep learning to tackle the sampling problem. Its follow-up, the denoising diffusion sampler (DDS) (Vargas et al., 2023a), takes a similar approach but adopts a variance preserving (VP) SDE rather than the variance exploding ones in Equation 5 and 7. We defer the derivation of VP modeling to Appendix B.2. The practical implementations of these prior works are performed in discrete time with constant step sizes, equivalent to the discrete formulation outlined above. Hence we also adopt the discrete-time formulation in the following sections for simplicity and better amenability to our proposed algorithm.
@@ 2.2 GFlowNets
Let $\mathcal{G}=(\mathcal{S},\mathcal{A})$ be a directed acyclic graph (DAG), where the set of vertices $\mathbf{s}\in\mathcal{S}$ are referred to as states, and the set of directed edges $(\mathbf{s}\to\mathbf{s}^{\prime})\in\mathcal{A}\subseteq\mathcal{S}\times\mathcal{S}$ represent the transitions between the states and are called actions. Given a designated initial state $\mathbf{s}_{0}$ with no incoming edges and a set of terminal states $s_{N}$ without outgoing edges, a complete trajectory can be defined as a sequence of states $\tau=(\mathbf{s}_{0}\to\mathbf{s}_{1}\to\ldots\to\mathbf{s}_{N})\in\mathcal{T}$. We define the forward policy $P_{F}(\mathbf{s}^{\prime}|\mathbf{s})$ as a distribution over the children of every non-terminal state $\mathbf{s}$, which induces a distribution over complete trajectories, $P_{F}(\tau)=\prod_{n=0}^{N-1}P_{F}(\mathbf{s}_{n+1}|\mathbf{s}_{n})$. This in turn defines a marginal distribution over terminating states $\mathbf{x}$ in the sense that $P_{T}(\mathbf{x})=\sum_{\tau\to\mathbf{x}}P_{F}(\tau)$, where we call $P_{T}(\cdot)$ the GFlowNet’s terminating distribution. The aim of GFlowNet learning is to obtain a policy such that $P_{T}(\cdot)\propto R(\cdot)$, where $R(\cdot):\mathcal{X}\to\mathbb{R}_{+}$ is a given reward function or unnormalized density, where we do not know the normalizing factor $Z=\sum_{\mathbf{x}}R(\mathbf{x})$. We also use the trajectory flow function $F(\tau)=ZP_{F}(\tau)$ to incorporate the effect of the partition function. Accordingly, the state flow function is the marginalization of the trajectory flow function $F(\mathbf{s})=\sum_{\tau\owns\mathbf{s}}F(\tau):\mathcal{S}\to\mathbb{R}_{+}$. Most previous GFlowNet works are about composite discrete objects such as molecules; on the other hand, Lahlou et al. (2023) provide theoretical support and initial numerical demonstration for GFlowNets on continuous and hybrid (continuous and discrete) spaces. Based on this, our work aims to further tackle more complex sampling problems in continuous domains, i.e. $\mathcal{S}\subseteq\mathbb{R}^{D}$.
@@ 3.2 Updating parameters with incomplete trajectories
This formulation is similar to the GFlowNet DB loss in Equation 9. Actually, it is not hard to see that the above modeling is a variant of GFlowNets. As Zhang et al. (2022a) has pointed out and as later adopted by Lahlou et al. (2023), there is a connection between diffusion modeling and GFlowNets. In the GFlowNet language, the target density $\mu(\cdot)$ is the terminal reward function, the temporal-spatial specification space $\{(n,\mathbf{x}_{n}):n\in\mathbb{N},\mathbf{x}_{n}\in\mathbb{R}^{D}\}$ is the GFlowNet state space $\mathcal{S}$, the forward transition probability $P_{F}(\cdot|\cdot)$ of the diffusion process $\mathcal{Q}$ is the GFlowNet forward policy, the backward transition probability $P_{B}(\cdot|\cdot)$ of the reference process $\mathcal{Q}^{\text{ref}}$ is the GFlowNet backward policy, and the helper network $F_{n}(\cdot)$ is the state flow $F(\mathbf{s})$. Notice here that the backward policy is fixed and does not contain learnable parameters. In our setting, we use the step index $n$ to distinguish between intermediate states $\{(n,\mathbf{x}_{n}):n<N,\mathbf{x}_{n}\in\mathbb{R}^{D}\}$ and terminating states $\{(N,\mathbf{x}_{N}),\mathbf{x}_{N}\in\mathbb{R}^{D}\}$ [^4] Without loss of generality, we slightly overuse the notations and simply use $\mathbf{x}_{n}$ to denote $(n,\mathbf{x}_{n})$.. What’s more, the goal of GFlowNets, which is to generate terminal states with probability proportional to the reward value, matches the goal of sampling in our setting.
@@
where $\lambda\in\mathbb{R}_{+}$ is a scalar for controlling assigned weights to different partial trajectories. For the modeling of the drift of the forward stochastic process, we use $\mathbf{f}(\cdot,n)/\sigma=\text{NN}_{1}(\cdot,n),+\text{NN}_{2}(n)\cdot\nabla\log\mu(\cdot)$, which includes two different neural networks and gradient information from unnormalized target densities in a similar form to Langevin dynamics (Zhang & Chen, 2022). We have conduct extensive ablation study about the parameterization, training objectives, exploration techniques, and other considerations, and display them in Appendix C.
@@
where $B\in\mathbb{N}_{+}$ is the number of particles (i.e., trajectories) used for the estimation. Similar to the importance weighted auto-encoders (Burda et al., 2016), our estimator provides a valid lower bound of the logarithm of the normalizing factor with any forward process $\mathcal{Q}$. In this work, we use the difference between estimated log partition function and the ground truth as an evaluation metric. See the Appendix C.2 for additional analysis. We also remark that this approach is similar to the neural importance sampling in Nicoli et al. (2020); Wirnsberger et al. (2020) in the sense that we use the forward process $\mathcal{Q}$ as the proposal distribution.
@@ 3.3 Discussion
Figure 2 shows that the variance of stochastic gradients with DGFS is much smaller than that with PIS; notice that the comparison is valid as we take the same neural network architecture for both algorithms. We hypothesize that DGFS benefits from the low variance behavior due to the following two reasons: (1) From the perspective of temporal difference learning (Sutton & Barto, 2018; Kearns & Singh, 2000), it has long been discussed that policy learning from shorter trajectory data would lower the variance at the cost of potentially higher bias. Before this work, it was impossible for sampling with stochastic process algorithms to update with partial trajectories, while in the proposed DGFS method, we use an additional flow network to make this feasible. In Figure 4 we also show that the flow network can be properly learned and thus does not introduce much bias, indicating that DGFS achieves a better bias-variance trade-off. (2) Probabilistic ML researchers have shown that for KL divergence based distribution learning, the gradient will not vanish even if the optimal distribution has been reached; see Roeder et al. (2017) and Section 4 of Xu et al. (2022). Later works including Vaitl et al. (2022a; b) propose a different path-gradient estimator to tackle this issue with a similar insight with Roeder et al. (2017); see the related work section. On the other hand, the DGFS objective does achieve zero gradient with the optimal solution (c.f. Appendix C.3), making it a more stable algorithm.
@@
Previous theoretical studies (Bortoli, 2022; Chen et al., 2022; Lee et al., 2022) have demonstrated that the distribution of terminal samples generated by a diffusion model can converge to the target distribution under mild assumptions if the control term is well learned. These findings are also applicable to DGFS since these proofs are independent of how the networks are trained. Additionally, Zhang et al. (2022a) have proven that a perfectly learned score term corresponds to a zero GFlowNet training loss under mild conditions, following theory of GFlowNets (Bengio et al., 2021; 2023). This provides assurance that a sufficiently well-trained DGFS can accurately sample from the target distribution.
@@ 4 Related works
MCMC methods (Brooks et al., 2011) are the most classical sampling algorithms, including Langevin dynamics (Welling & Teh, 2011), Hamiltonian Monte Carlo (Duane et al., 1987; Hoffman & Gelman, 2011; Chen et al., 2014), and others (Huang et al., 2023). With MCMC operators as building blocks, sequential MC methods (Liu, 2001; Doucet et al., 2001) generate samples ancestrally through a series of annealed targets. On the other hand, people turn to parametric models such as deep networks to learn samplers via variational inference (Rezende & Mohamed, 2015). One the most commonly adopted models is the normalizing flow (Dinh et al., 2014) due to its tractability and expressiveness. Various work have been conducted to improve normalizing flow based sampling (Müller et al., 2018; Gao et al., 2020; Wu et al., 2020; Chen et al., 2020; Gabri’e et al., 2021; Arbel et al., 2021; Midgley et al., 2022; de G. Matthews et al., 2022; Caselle et al., 2022; Gao et al., 2023). These methods support asymptotically unbiased estimation as shown in Nicoli et al. (2020). What’s more, Vaitl et al. (2022a) propose to to improve the gradient estimation of normalizing flow parameters with the path-gradient technique; Vaitl et al. (2022b) improve continuous normalizing flows based methods (Chen et al., 2018) with a similar technique. These methods have wide applications in natural sciences, for example in lattice field theories (Nicoli et al., 2021; 2023) and targeted free energy estimation (Wirnsberger et al., 2020).
Recently, many researchers have developed powerful diffusion process to model distributions in the generative modeling community (Vincent, 2011; Sohl-Dickstein et al., 2015; Ho et al., 2020a; Song et al., 2020; Du et al., 2022). This inspires people to study how to utilize diffusion process as samplers to learn distribution from given unnormalized target densities (Tzen & Raginsky, 2019; Zhang & Chen, 2022; Vargas et al., 2021; Geffner & Domke, 2022; Doucet et al., 2022; Vargas et al., 2023a; Richter et al., 2023; Vargas & Nusken, 2023; Vargas et al., 2023c; b). Our work also falls into this category. Specifically, Vargas et al. (2023a) can be seen as a special case of the method proposed in Berner et al. (2022).
@@
Table 1: Experimental results on benchmarking target densities. For each algorithm, we report its absolute estimation bias for the log normalizing factor and the corresponding standard deviation. Our method achieves the best performance among the diffusion modeling-based samplers.
@@
FAB w/ Buffer [^5] We used the fab-jax results. The FAB method utilizes the annealed importance sampling computation during the training period. $0.003$±$\pm 0.0005$ $0.0022$±$\pm 0.0005$ $0.032$±$\pm 0.004$ N/A $0.19$±$\pm 0.04$
@@
DDS $0.028$±$\pm 0.013$ $0.416$±$\pm 0.094$ $1.154$±$\pm 0.626$ $1.740$±$\pm 1.158$ N/A [^6] Due to time constraints, we are unable to reproduce DDS performance using PyTorch for this task. Instead, we encountered “not a number”, probably due to a mismatch in implementation details or other unspecified difference between PyTorch and Jax. The DDS results reported in its original paper can be found here.
DGFS $0.019$±$\pm 0.008$ $0.274$±$\pm 0.014$ $0.904$±$\pm 0.067$ $0.180$±$\pm 0.083$ $8.974$±$\pm 1.169$
GFlowNet (Bengio et al., 2023) is a family of generalized variational inference algorithms with a focus on treating the sampling process as a sequential decision-making process. It is original proposed to sample diversity-seeking objects in structured scientific domains such as biological seuqence and molecule design (Bengio et al., 2021; Jain et al., 2022; 2023b; 2023a; Shen et al., 2023; Zhang et al., 2023c; Hernández-García et al., 2023; Kim et al., 2023). However, people have developed GFlowNet theories and methodologies to be suitable for generating general objects with composite structures. GFlowNet has rich connection with existing probabilistic ML methods (Zhang et al., 2022b; Zimmermann et al., 2022; Malkin et al., 2023; Zhang et al., 2022a; Hu et al., 2023; Ma et al., ). Malkin et al. (2023) also show that GFlowNet has the capability of training with off-policy data compared to previous amortized inference methods. As a policy learning framework under Markov decision processes (MDPs), researchers have also drawn fruitful connections between GFlowNet and reinforcement learning (Pan et al., 2023b; c; a). GFlowNet has also been demonstrated to have wide application in causal discovery (Deleu et al., 2022; 2023; Atanackovic et al., 2023), feature attribution (Li et al., 2023a), network structure learning (Liu et al., 2022), job scheduling (Zhang et al., 2023a), and graph combinatorial optimization (Zhang et al., 2023b). Most prior works focus on structured discrete data space. Lahlou et al. (2023) first provides a sound theoretical framework and initial numerical demonstration for using GFlowNet on continuous space. Our work follows this framework and scales up the performance. A prior study (Li et al., 2023b) initially explored GFlowNets in a continuous space, whereas their theory is demonstrated to have non-trivial flaws when considering integration under continuous measure as indicated by Lahlou et al. (2023).